OpenStudy (anonymous):

If x, y, and z represent three numbers such that y=z-4 and z=x-8, what is the result when the median of the numbers is subtracted from the average of the numbers?

OpenStudy (anonymous):

\[\begin{cases} x\\ y=z-4\\ z=x-8 \end{cases}~\Rightarrow~\color{red}{\begin{cases}x=x\\y=(x-8)-4\\z=x-8\end{cases}}\] The average of x, y, and z is \[average=\frac{x+y+z}{3}=\frac{x+(x-12)+(x-8)}{3}=\frac{3x-20}{3}\] The median number is the "middle" number. So if you have \(x<y<z\), for example, then \(y\) is the median number. In this case, judging by the red set of equations, \(z\) is the median number. This is because \[x-12<x-8<x~\iff~y<z<x\] Now find the difference between the median and average.

OpenStudy (anonymous):

Would you just add the 12 and the 8 on all sides?

OpenStudy (loser66):

thanks @SithsAndGiggles

OpenStudy (anonymous):

No problem @Loser66. @sakigirl, what do you mean by "all sides?" The question wants you to find \[average-median\] which we've revealed to be equivalent to \[\frac{3x-20}{3}-(x-8)\]

OpenStudy (anonymous):

And so to subtract, would you multiply, 1/3 to x-8?

OpenStudy (anonymous):

Not quite. You'd be missing a factor of 3 in the numerator. You have to multiply by 3/3: \[\frac{3x-20}{3}-\frac{3(x-8)}{3}\]

OpenStudy (anonymous):

So would it be -24/3 which is -8?

OpenStudy (anonymous):

@SithsAndGiggles Sorry, you have no clue how incredibly dumb I feel, haha

OpenStudy (anonymous):

Close! I think you made a small mistake with adding: \[\frac{3x-20}{3}-\frac{3(x-8)}{3}\\ \frac{3x-20}{3}-\frac{3x-24}{3}\\ \frac{(3x-20)-(3x-24)}{3}\\ \frac{3x-20-3x+24}{3}\\ \frac{-20+24}{3}\\ ~~~~~~~~~\vdots\]

OpenStudy (anonymous):

@SithsAndGiggles Thank you :)

OpenStudy (anonymous):

You're welcome!